# Transferability of Adversarial Attacks in Video-based MLLMs: A Cross-modal Image-to-Video Approach

### v3

@@ 4.3 Ablation study
Different perturbation propagation methods may yield varying results. We set the key-frame ratio $\beta=30\%$ and test three distinct perturbation propagation methods: Direct Propagation (DP), Optical Flow-based Propagation (dosovitskiy2015flownet) (OFP), and Bidirectional Linear Interpolation Propagation (dai2017deformable) (BP). The details of propagation methods are in Appendix B.1. As shown in Tab. 4, DP achieves the most significant improvement in AASR. Due to the high similarity between consecutive frames, DP proves to be a simple yet effective method. OFP suffers from added complexity and may distort the perturbation, resulting in lower effectiveness. BP’s slightly lower performance stems from its reliance on interpolation, which may dilute perturbation intensity compared to DP’s direct application. Therefore, we adopt DP in I2V-MLLM.
Influence of different projectors. The projectors of I-MLLMs are typically either FC-Linear (LLaVA (liu2023visualinstructiontuning), mPLUG-Owl (ye2023mplug)) or Q-Former (BLIP-2, InstructBLIP, MiniGPT-4). FC-Linear maps the visual features extracted by the vision encoder into the latent space of the LLM, whereas Q-Former further aligns visual and textual features before passing them to the LLM, enabling richer multimodal interactions. When Q-Former-based I-MLLM is employed as the surrogate, the projector attack module’s loss function is $\mathcal{L}_{P}=\mathcal{L}_{P_{v}}+\mathcal{L}_{P_{v2t}}$, where as for FC-Linear-based I-MLLM, it becomes $\mathcal{L}_{P}=\mathcal{L}_{P_{v}}$. As shown in Tab. 5, irrespective of the choice of surrogate model and projector, the adversarial transferability of I2V-MLLM significantly surpasses that of the vanilla attack. I-MLLMs equipped with Q-Former consistently outperform those with FC-Linear in generating transferable adversarial samples. This underscore the critical role of multimodal interactions in adversarial attacks on V-MLLMs. Therefore, in our main experiments, we exclusively employ Q-Former-based I-MLLMs as surrogate models to fully demonstrate the effectiveness of the I2V-MLLM approach.
@@ Acknowledgments
Transferability of Adversarial Attacks in Video-based MLLMs:
@@ Appendix A Motivation
Lacking generalization in perturbing video features. The gains from perturbation propagation are limited due to variations in how V-MLLMs extract video features, which often causes the perturbations to overfit to the features of the surrogate model’s video encoder, which reduces the generalizability of perturbations. Enhancing transferability requires targeting common elements across these features. Thus, we focus on lower-level image features. The I2V attack (wei2021crossmodaltransferableadversarialattacks), which perturbs each video frame to disrupt image features, demonstrates that using image models as surrogates can effectively generate adversarial samples for video models. Experimental results in row 3, 4 and 5 of Tab. 1 in the main content confirm that targeting image features in video frames significantly improves the transferablity of adversarial video samples.
@@ B.1 Experiment setting
Models. We perform our proposed approach on five I-MLLMs: BLIP-2 (li2023blip2bootstrappinglanguageimagepretraining), InstructBLIP (dai2023instructblipgeneralpurposevisionlanguagemodels), MiniGPT-4 (zhu2023minigpt4enhancingvisionlanguageunderstanding), LLaVA (liu2023visualinstructiontuning) and mPLUG-owl (ye2023mplug). Our proposed methods are evaluated on four different V-MLLMs: Chat-UniVi (jin2024chatuniviunifiedvisualrepresentation), LLaVA-Next-Video (zhang2024videoinstructiontuningsynthetic), VideoChat (li2024videochatchatcentricvideounderstanding) and Video-LLaMA (zhang2023videollamainstructiontunedaudiovisuallanguage), each with a Vicuna-7B (vicuna2023) as the LLM.
@@
Propagation methods. Direct Propagation (DP), Optical Flow-based Propagation (dosovitskiy2015flownet) (OFP), and Bidirectional Linear Interpolation Propagation (dai2017deformable) (BP) are three different perturbation propagation methods. In DP, the adversarial perturbation is directly applied to each corresponding frame segment without any adjustments, simply propagating the perturbation along the frame sequence segments. OFP calculates the optical flow changes between adjacent frames and dynamically adjusts and propagates the adversarial perturbation based on the motion information between frames, facilitating a perturbation transfer that better aligns with frame-to-frame motion patterns. BP employs bidirectional linear interpolation to propagate perturbations from adjacent key frames to the intermediate frames, thereby smoothly transmitting adversarial perturbations across the entire frame sequence.
@@
Metrics. We use Attack Success Rate (ASR) to evaluate the effectiveness of adversarial examples on VideoQA tasks. It measures the percentage of successful attacks on questions the model answered correctly for clean videos. Answer correctness is evaluated using GPT-4o-mini (openai2023gpt4omini), which checks whether the model’s prediction semantically aligns with the ground truth. We also provide the average ASR (AASR) across all evaluated V-MLLMs. A higher ASR or AASR indicates better adversarial transferability. To evaluate the model’s overall performance when encountering adversarial videos, we further employ GPT-assisted methods (maaz2024videochatgptdetailedvideounderstanding) to assess Accuracy (Acc.) and GPT-Score. We use GPT-4o-mini (openai2023gpt4omini) as the evaluation model due to its superior performance in text understanding and its cost efficiency. Fig. 5 illustrates an example of evaluating the Zero-Shot VideoQA task with GPT-4o-mini. An adversarial attack is considered successful if the adversarial video sample causes the V-MLLM to change its response to a question from correct (Accuracy: 1) to incorrect (Accuracy: 0).
@@
Figure 5: An example of using GPT-4o-mini to evaluate Accuracy and GPT Score for the VideoQA task, following the methodology in (maaz2024videochatgptdetailedvideounderstanding).
@@
Since traditional video captioning metrics (Wang_2018_CVPR) such as BLEU, CIDEr, SPICE are not well-suited for evaluating the detailed video captions generated by V-MLLM, we employ GPT-4o-mini as the evaluation model and use the Captioning Score to assess V-MLLM’s performance on the video captioning task. The Captioning Score measures the quality of the model’s predictions by assigning a relative score on a scale from 0 to 100 (see Fig. 6). A higher Captioning Score indicates that the generated captions are closer to the reference captions.
@@ B.2 Cross-task transferability
Multimodal interactions and alignments are the core components of multimodal learning, regardless of the specific task. Therefore, we conduct extensive experiments to investigate the cross-task transferability of I2V-MLLM from VideoQA to video captioning tasks. Since traditional video captioning metrics (Wang_2018_CVPR) such as BLEU, CIDEr, SPICE are not well-suited for evaluating the detailed video captions generated by V-MLLMs, we employ GPT-4o-mini as the evaluation model and use the Captioning Score to assess V-MLLM’s performance on the video captioning task. The Captioning Score measures the quality of the model’s predictions by assigning a relative score on a scale from 0 to 100. A higher Captioning Score indicates that the generated captions are closer to the reference captions. We directly transfer the adversarial samples generated from the VideoQA on the MSVD-QA dataset to the video captioning tasks on the MSVD dataset. As shown in Tab. 7, I2V-MLLM causes the largest decrease in the Captioning Scores across all target models, achieving the best attack performance, even surpassing Vanilla and FMM attack methods in the white-box setting. The white-box Vanilla attack, which achieves optimal performance in the VideoQA tasks, performs sub-optimally here because it is specifically designed for the VideoQA task, resulting in slightly weaker cross-task transferability. This highlights that I2V-MLLM not only exhibits cross-model transferability but also demonstrates strong cross-task transferability.
@@ B.3 Collaboration with textual attacks
In real-world scenarios, attackers often employ a wide range of strategies to attack V-MLLMs, potentially introducing both adversarial video samples and adversarial text into the target models simultaneously. To account for this more practical and comprehensive attack scenario, we conducted additional experiments. Specifically, we applied textual attacks on the MSVD-QA dataset using the Bert-Attack method (li2020bert), modifying only a single character in each question to maximally disrupt the semantic alignment between the video and text.
@@
Table 8: Attack success rates (ASR, %) on the MSVD-QA validation set for Zero-Shot VideoQA tasks. Bert-Attack from (li2020bert).
@@ B.4 Ablation study
Influence of perturbation bound. To further investigate the impact of perturbation magnitude on attack performance, we tested the ASR under different perturbation bounds ($\epsilon$), including 2, 4, 8, 16, and 32. As shown Tab. 11, the ASR steadily increases with higher perturbation bounds, reaching its peak when the perturbation bound is set to 32. As illustrated in Figs. 9 and 10, with the increase in perturbation bound, the perturbations become more pronounced. To balance attack performance and video vividness, we selected a perturbation bound of 16, which is also a common choice in traditional video attack methods (wei2021crossmodaltransferableadversarialattacks; kim2023breakingtemporalconsistencygenerating; wang2023global).
@@
Multi-faceted video understanding tasks assess whether V-MLLMs have comprehended the content of a video by posing a range of questions about it. Following Maaz et al. (maaz2024videochatgptdetailedvideounderstanding), we use a subset of the ActivityNet-200 (Heilbron_Escorcia_Ghanem_Niebles_2015) dataset and employ GPT-4o-mini to evaluate the model’s responses to adversarial examples from five perspectives: Correctness, Detail Orientation, Contextual Understanding, Temporal Understanding, and Consistency. We compare our proposed I2V-MLLM attack with the Vanilla attack on four V-MLLMs, using clean samples as a reference. Evaluations are performed on white-box Vanilla attack and I2V-MLLM attack (using BLIP-2 as a surrogate model).
@@ B.6 Analysis
Discussion. To experimentally validate the effectiveness of Eq (6) in the main content, we analyze how the cosine similarity between adversarial and benign features in I-MLLM/V-MLLMs evolves as the iteration number increases. The Pearson Correlation Coefficient (PCC) (Schmee_1986) is used to quantify the linear correlation between cosine similarity trends computed from both I-MLLM and V-MLLMs. Fig. 11 presents the PCC analysis of these trends, using BLIP-2 and four different V-MLLMs. As shown, all PCC values exceed 0.90, indicating a strong positive linear relationship between the directional changes of intermediate features in I-MLLM and V-MLLM. This suggests that perturbations in I-MLLM’s image features can effectively disrupt the intermediate features of video samples in V-MLLMs. Notably, the PCC values between BLIP-2 and VideoChat, as well as Video-LLaMA, are exactly 1, which aligns with the highest ASR values observed for these models in Tab. 2 in the main content. The slightly lower PCC values with Chat-UniVi and LLaVA-NeXT-Video correspond to the lower ASR values, demonstrating that a higher PCC between I-MLLM and V-MLLMs indicates better adversarial transferability.

### v4

@@ 4.3 Ablation study
Different perturbation propagation methods may yield varying results. We set the key-frame ratio $\beta=30\%$ and test three distinct perturbation propagation methods: Direct Propagation (DP), Optical Flow-based Propagation (Dosovitskiy et al. 2015) (OFP), and Bidirectional Linear Interpolation Propagation (Dai et al. 2017) (BP). The details of propagation methods are in Appendix B.1. As shown in Tab. 4, DP achieves the most significant improvement in AASR. Due to the high similarity between consecutive frames, DP proves to be a simple yet effective method. OFP suffers from added complexity and may distort the perturbation, resulting in lower effectiveness. BP’s slightly lower performance stems from its reliance on interpolation, which may dilute perturbation intensity compared to DP’s direct application. Therefore, we adopt DP in I2V-MLLM.
Influence of different projectors. The projectors of I-MLLMs are typically either FC-Linear (LLaVA (Liu et al. 2023), mPLUG-Owl (Ye et al. 2023)) or Q-Former (BLIP-2, InstructBLIP, MiniGPT-4). FC-Linear maps the visual features extracted by the vision encoder into the latent space of the LLM, whereas Q-Former further aligns visual and textual features before passing them to the LLM, enabling richer multimodal interactions. When Q-Former-based I-MLLM is employed as the surrogate, the projector attack module’s loss function is $\mathcal{L}_{P}=\mathcal{L}_{P_{v}}+\mathcal{L}_{P_{v2t}}$, where as for FC-Linear-based I-MLLM, it becomes $\mathcal{L}_{P}=\mathcal{L}_{P_{v}}$. As shown in Tab. 5, irrespective of the choice of surrogate model and projector, the adversarial transferability of I2V-MLLM significantly surpasses that of the vanilla attack. I-MLLMs equipped with Q-Former consistently outperform those with FC-Linear in generating transferable adversarial samples. This underscore the critical role of multimodal interactions in adversarial attacks on V-MLLMs. Therefore, in our main experiments, we exclusively employ Q-Former-based I-MLLMs as surrogate models to fully demonstrate the effectiveness of the I2V-MLLM approach.
@@ Appendix A Motivation
Lacking generalization in perturbing video features. The gains from perturbation propagation are limited due to variations in how V-MLLMs extract video features, which often causes the perturbations to overfit to the features of the surrogate model’s video encoder, which reduces the generalizability of perturbations. Enhancing transferability requires targeting common elements across these features. Thus, we focus on lower-level image features. The I2V attack (Wei et al. 2021), which perturbs each video frame to disrupt image features, demonstrates that using image models as surrogates can effectively generate adversarial samples for video models. Experimental results in row 3, 4 and 5 of Tab. 1 in the main content confirm that targeting image features in video frames significantly improves the transferablity of adversarial video samples.
@@ B.1 Experiment setting
Models. We perform our proposed approach on five I-MLLMs: BLIP-2 (Li et al. 2023), InstructBLIP (Dai et al. 2023), MiniGPT-4 (Zhu et al. 2023), LLaVA (Liu et al. 2023) and mPLUG-owl (Ye et al. 2023). Our proposed methods are evaluated on four different V-MLLMs: Chat-UniVi (Jin et al. 2024), LLaVA-Next-Video (Zhang et al. 2024b), VideoChat (Li et al. 2024b) and Video-LLaMA (Zhang et al. 2023), each with a Vicuna-7B (Chiang et al. 2023) as the LLM.
@@
Propagation methods. Direct Propagation (DP), Optical Flow-based Propagation (Dosovitskiy et al. 2015) (OFP), and Bidirectional Linear Interpolation Propagation (Dai et al. 2017) (BP) are three different perturbation propagation methods. In DP, the adversarial perturbation is directly applied to each corresponding frame segment without any adjustments, simply propagating the perturbation along the frame sequence segments. OFP calculates the optical flow changes between adjacent frames and dynamically adjusts and propagates the adversarial perturbation based on the motion information between frames, facilitating a perturbation transfer that better aligns with frame-to-frame motion patterns. BP employs bidirectional linear interpolation to propagate perturbations from adjacent key frames to the intermediate frames, thereby smoothly transmitting adversarial perturbations across the entire frame sequence.
@@
Metrics. We use Attack Success Rate (ASR) to evaluate the effectiveness of adversarial examples on VideoQA tasks. It measures the percentage of successful attacks on questions the model answered correctly for clean videos. Answer correctness is evaluated using GPT-4o-mini (OpenAI 2024a), which checks whether the model’s prediction semantically aligns with the ground truth. We also provide the average ASR (AASR) across all evaluated V-MLLMs. A higher ASR or AASR indicates better adversarial transferability. To evaluate the model’s overall performance when encountering adversarial videos, we further employ GPT-assisted methods (Maaz et al. 2024) to assess Accuracy (Acc.) and GPT-Score. We use GPT-4o-mini (OpenAI 2024a) as the evaluation model due to its superior performance in text understanding and its cost efficiency. Fig. 5 illustrates an example of evaluating the Zero-Shot VideoQA task with GPT-4o-mini. An adversarial attack is considered successful if the adversarial video sample causes the V-MLLM to change its response to a question from correct (Accuracy: 1) to incorrect (Accuracy: 0).
@@
Figure 5: An example of using GPT-4o-mini to evaluate Accuracy and GPT Score for the VideoQA task, following the methodology in (Maaz et al. 2024).
@@
Since traditional video captioning metrics (Wang et al. 2018) such as BLEU, CIDEr, SPICE are not well-suited for evaluating the detailed video captions generated by V-MLLM, we employ GPT-4o-mini as the evaluation model and use the Captioning Score to assess V-MLLM’s performance on the video captioning task. The Captioning Score measures the quality of the model’s predictions by assigning a relative score on a scale from 0 to 100 (see Fig. 6). A higher Captioning Score indicates that the generated captions are closer to the reference captions.
@@ B.2 Cross-task transferability
Multimodal interactions and alignments are the core components of multimodal learning, regardless of the specific task. Therefore, we conduct extensive experiments to investigate the cross-task transferability of I2V-MLLM from VideoQA to video captioning tasks. Since traditional video captioning metrics (Wang et al. 2018) such as BLEU, CIDEr, SPICE are not well-suited for evaluating the detailed video captions generated by V-MLLMs, we employ GPT-4o-mini as the evaluation model and use the Captioning Score to assess V-MLLM’s performance on the video captioning task. The Captioning Score measures the quality of the model’s predictions by assigning a relative score on a scale from 0 to 100. A higher Captioning Score indicates that the generated captions are closer to the reference captions. We directly transfer the adversarial samples generated from the VideoQA on the MSVD-QA dataset to the video captioning tasks on the MSVD dataset. As shown in Tab. 7, I2V-MLLM causes the largest decrease in the Captioning Scores across all target models, achieving the best attack performance, even surpassing Vanilla and FMM attack methods in the white-box setting. The white-box Vanilla attack, which achieves optimal performance in the VideoQA tasks, performs sub-optimally here because it is specifically designed for the VideoQA task, resulting in slightly weaker cross-task transferability. This highlights that I2V-MLLM not only exhibits cross-model transferability but also demonstrates strong cross-task transferability.
@@ B.3 Collaboration with textual attacks
In real-world scenarios, attackers often employ a wide range of strategies to attack V-MLLMs, potentially introducing both adversarial video samples and adversarial text into the target models simultaneously. To account for this more practical and comprehensive attack scenario, we conducted additional experiments. Specifically, we applied textual attacks on the MSVD-QA dataset using the Bert-Attack method (Li et al. 2020), modifying only a single character in each question to maximally disrupt the semantic alignment between the video and text.
@@
Table 8: Attack success rates (ASR, %) on the MSVD-QA validation set for Zero-Shot VideoQA tasks. Bert-Attack from (Li et al. 2020).
@@ B.4 Ablation study
Influence of perturbation bound. To further investigate the impact of perturbation magnitude on attack performance, we tested the ASR under different perturbation bounds ($\epsilon$), including 2, 4, 8, 16, and 32. As shown Tab. 11, the ASR steadily increases with higher perturbation bounds, reaching its peak when the perturbation bound is set to 32. As illustrated in Figs. 9 and 10, with the increase in perturbation bound, the perturbations become more pronounced. To balance attack performance and video vividness, we selected a perturbation bound of 16, which is also a common choice in traditional video attack methods (Wei et al. 2021; Kim et al. 2023; Wang et al. 2023).
@@
Multi-faceted video understanding tasks assess whether V-MLLMs have comprehended the content of a video by posing a range of questions about it. Following Maaz et al. (Maaz et al. 2024), we use a subset of the ActivityNet-200 (Heilbron et al. 2015) dataset and employ GPT-4o-mini to evaluate the model’s responses to adversarial examples from five perspectives: Correctness, Detail Orientation, Contextual Understanding, Temporal Understanding, and Consistency. We compare our proposed I2V-MLLM attack with the Vanilla attack on four V-MLLMs, using clean samples as a reference. Evaluations are performed on white-box Vanilla attack and I2V-MLLM attack (using BLIP-2 as a surrogate model).
@@ B.6 Analysis
Discussion. To experimentally validate the effectiveness of Eq (6) in the main content, we analyze how the cosine similarity between adversarial and benign features in I-MLLM/V-MLLMs evolves as the iteration number increases. The Pearson Correlation Coefficient (PCC) (Schmee 1986) is used to quantify the linear correlation between cosine similarity trends computed from both I-MLLM and V-MLLMs. Fig. 11 presents the PCC analysis of these trends, using BLIP-2 and four different V-MLLMs. As shown, all PCC values exceed 0.90, indicating a strong positive linear relationship between the directional changes of intermediate features in I-MLLM and V-MLLM. This suggests that perturbations in I-MLLM’s image features can effectively disrupt the intermediate features of video samples in V-MLLMs. Notably, the PCC values between BLIP-2 and VideoChat, as well as Video-LLaMA, are exactly 1, which aligns with the highest ASR values observed for these models in Tab. 2 in the main content. The slightly lower PCC values with Chat-UniVi and LLaVA-NeXT-Video correspond to the lower ASR values, demonstrating that a higher PCC between I-MLLM and V-MLLMs indicates better adversarial transferability.